# Input Convex Lipschitz RNN: A Fast and Robust Approach for Engineering Tasks

### v3

@@ V Limitations and Future Works
Both ICRNN and our ICLRNN exhibit limitations in tasks requiring the modeling of very long time-dependent sequences, such as the byte-level text classification task, where both achieve an accuracy of approximately 50% in this binary sentiment classification task. This limitation stems from the challenge of the exploding gradient problem and lack of representative power due to non-negative weights, which both methods encounter when attempting to model extremely lengthy sequences, and we provide a theoretical explanation as follows.
While the Lipschitz constraint may alleviate the exploding gradient problem by bounding the weights, integrating input convexity introduces a new challenge by raising this upper bound. Applying the Lipschitz constraint first to reduce the spectral norm to 1, followed by imposing the input convex constraint, yields a neural network no longer 1-Lipschitz constrained but rather $n\sqrt{r(W)}$-constrained, where $n$ represents the number of time-steps. Consequently, in modeling very long time-dependent sequences, the Input Convex structure exacerbates the exploding gradient problem, rendering it unmitigated even with Layer Normalization. For readers interested in tackling the challenges of the exploding and vanishing gradient problem within RNNs, we recommend exploring several noteworthy and comprehensive works [33, 34, 35].
@@
We acknowledge this limitation as inherent to our proposed method, thereby constraining its applicability. However, despite this drawback, we opt to retain the Input Convex structure, as it demonstrates efficacy in many engineering applications where the length of time-series data is not as long as those in Long Range Arena benchmark. For example, in the first case study where we develop a model for real-world solar irradiance prediction, the prediction horizon is relatively short (e.g., an hour or a few hours ahead) since this model is used for real-time energy planning. In the second example where we develop a model for real-time optimization of CSTR, again, the model is to predict one sampling period only, since the controller is typically implemented in a feedback manner with the true measurements of states received at every time step. Therefore, there are many applications in engineering disciplines that focus on short-term prediction, where the proposed ICLRNN will play an important role due to its superiority in computational efficiency and robustness compared to traditional RNNs.
Furthermore, because of the input convex architecture, all ICNNs excel in regression tasks rather than classification tasks, including ICLRNN. This is because, to maintain input convexity, all activation functions must be convex, prohibiting the use of certain activation functions such as Sigmoid, which is standard for classification tasks. However, given our focus in this work on developing a fast and robust neural network tailored for optimization tasks, which are regression-based, we can disregard this limitation.
VI Conclusion
@@
In conclusion, this work introduces an Input Convex Lipschitz RNN to improve computational efficiency and non-adversarial robustness of neural networks in various engineering applications such as real-world solar irradiance prediction, and process modeling and optimization. Theoretical analyses of its input convexity and Lipschitz continuity properties are provided. Furthermore, from our experiments, we observed that both LRNN and ICRNN require significant computational resources. LRNN hinders the performance of optimizer. Additionally, ICRNN encounters challenges related to exploding gradient problems when operating within large hypothesis space sizes and proved sensitive to input perturbations. Despite potential limitations, the ICLRNN model was demonstrated to surpass state-of-the-art recurrent units across various engineering tasks, ranging from simulated studies to real-world applications, in terms of computational efficiency and non-adversarial robustness.
@@ Appendix A Proofs
Definition 1 ([36]).
A function $f:X\rightarrow\mathbb{R}$ is convex if and only if the following inequality holds $\forall\leavevmode\nobreak\ (x_{1},x_{2})\in X$, with $x_{1}\neq x_{2}$, and $\forall\leavevmode\nobreak\ \lambda\in(0,1)$:
@@
Definition 2 ([37]).
A function $f:X\rightarrow\mathbb{R}$ is Lipschitz continuous with Lipschitz constant $L$ (i.e., L-Lipschitz) if and only if the following inequality holds $\forall\leavevmode\nobreak\ (x_{1},x_{2})\in X$:
@@
The following lemmas are provided to prove Proposition 2/4 regarding the Lipschitz continuity of ICLRNN.
Lemma 1 ([37]).
@@
$\begin{split}L(f)\leq&L(g_{t})\times L(W_{t}^{y})\times L(g_{t-1})\times(L(W_{t-1}^{x})\\ &+L(W_{t-1}^{h})\times...L(g_{1})\times(L(W_{1}^{x})+L(W_{1}^{h})))\\ \end{split}$
If we ensure that the Lipschitz constants of all $W_{i}$ in ICLRNN are $\mathcal{O}(1)$, the ICLRNN for classification tasks with ReLU activation for hidden states and Softmax activation for output layer has a Lipschitz constant of $\mathcal{O}(1)$, while the ICLRNN for regression tasks with ReLU activation for hidden states and Linear activation for output layer has a Lipschitz constant of $\mathcal{O}(1)$ (i.e., Lemma 3 and Lemma 4). ∎
The following lemmas are provided to prove Proposition 3/5 regarding the convexity of ICLRNN.
Lemma 5 ([36]).
@@
Proof.
By unrolling the recurrent operations over time $t$ in Eq. (1), the output of ICLRNN can be described as a function of Eq. (5). The output of an ICLRNN $y=f(x)$ is convex with respect to $x$ (i.e., Lemma 5 and Lemma 6). ∎
@@
Appendix B Real-World Solar Irradiance Prediction Results
@@
Fig. 5 illustrates the production pipeline adopted by LHT Holdings. Meanwhile, in Fig. 6, a depiction of the different energy sources employed within the production process is presented. Fig. 7 showcases the physical setup of the Solar PV system, encompassing solar panels, inverters, batteries, meters, humidity and wind sensors, module temperature sensors, and irradiance sensors. Data is systematically recorded on a minute-by-minute basis and seamlessly uploaded online into the system.
@@
Figure 6: LHT Holdings energy sources for manufacturing.
Figure 7: LHT Holdings Solar PV system.
@@
Moving to the bottom two figures, the LRNN and ICLRNN prediction results for the same date, December 28, 2023, are presented. The LRNN, constrained by Lipschitz continuity, exhibits significantly improved performance compared to the RNN and LSTM models. However, it still encounters challenges in accurately predicting solar irradiance during sudden fluctuations, and its excessively high FLOPs signify its computational intensity (i.e., 2,498,562 for LRNN compared to 399,362 for ICLRNN). Moreover, slight fluctuations in the LRNN predictions occur during periods of minimal solar irradiance at the start and end of the day. Notably, these observations persist across all trials and are not specific solely to December 28, 2023. It is clearly shown in Fig. 2 that our ICLRNN model surpasses state-of-the-art models, demonstrating superior predictive capabilities across various scenarios.
@@
Appendix C MPC-Based CSTR Results
Fig. 3 showcases the model’s testing accuracy with respect to varying degrees of Gaussian noise applied to the input. These findings indicate ICLRNN’s non-adversarial robustness across diverse hypothesis space sizes with respect to different Gaussian noise levels. Notably, ICLRNN exhibits superior computational efficiency, as evidenced by its notably lower number of FLOPs compared to other models.
While ICRNN performs well against various degrees of Gaussian noise within smaller hypothesis spaces, it encounters challenges when the hypothesis space expands, succumbing to the issue of the exploding gradient problem, as evidenced in previous task. LRNN demonstrates comparable performance to ICLRNN specifically with a layer size of (512, 512). However, it is computationally demanding, characterized by substantially higher FLOPs (i.e., 9,992,212 for LRNN compared to 1,599,508 for ICLRNN). This computational intensity stands as a significant drawback despite its competitive performance.
We now delve into the control performance assessment of NN-based LMPC. To assess and compare the convergence rates of various NN-based LMPC, we conduct different NN-based LMPC under a fixed timeframe and repeat with different initial conditions. Figs. 9 and 10 provide compelling evidence that the ICLRNN-based LMPC achieves convergence, unlike other NN-based LMPC, within a fixed timeframe. This outcome distinctly indicates that ICLRNN-based LMPC exhibits the fastest convergence speed under identical settings.
In Fig. 9, the left panels show the results corresponding to the initial condition set at $[-1.5\leavevmode\nobreak\ kmol/m^{3},70\leavevmode\nobreak\ K]$. Specifically, the top left panel displays the concentration profile, the middle left panel shows the temperature, and the bottom left panel exhibits the converging path. Additionally, the right panels showcase the results from the initial condition at $[1.5\leavevmode\nobreak\ kmol/m^{3},-70\leavevmode\nobreak\ K]$, featuring the concentration profile in the top right panel, temperature profile in the middle right panel, and the converging path in the bottom right panel. Moving to Fig. 10, the left panels depict the results obtained from the initial condition of $[-1.25\leavevmode\nobreak\ kmol/m^{3},50\leavevmode\nobreak\ K]$. Here, the top left panel represents the concentration, the middle left panel illustrates the temperature, and the bottom left panel shows the converging path. Similarly, the right panels exhibit the results of the initial condition at $[1.25\leavevmode\nobreak\ kmol/m^{3},-50\leavevmode\nobreak\ K]$.
Figure 9: Concentration profile (top left), temperature profile (middle left), converging path (bottom left) in a fixed timeframe with an initial condition at $[-1.5\leavevmode\nobreak\ kmol/m^{3},70\leavevmode\nobreak\ K]$ under LMPC, and concentration profile (top right), temperature profile (middle right), converging path (bottom right) in a fixed timeframe with an initial condition at $[1.5\leavevmode\nobreak\ kmol/m^{3},-70\leavevmode\nobreak\ K]$.
Figure 10: Concentration profile (top left), temperature profile (middle left), converging path (bottom left) in a fixed timeframe with an initial condition at $[-1.25\leavevmode\nobreak\ kmol/m^{3},50\leavevmode\nobreak\ K]$, and concentration profile (top right), temperature profile (middle right), converging path (bottom right) in a fixed timeframe with an initial condition at $[1.25\leavevmode\nobreak\ kmol/m^{3},-50\leavevmode\nobreak\ K]$.

### v4

@@ 5 More applications and discussions
During our experiments, we observed that ICRNN faces challenges related to the exploding gradient problem, prompting us to discuss this issue and its relevance to the exploding and vanishing gradients phenomenon. Fortunately, as a beneficial side-effect of addressing our aforementioned objectives, our proposed method exhibits reduced susceptibility to the exploding gradient problem to a certain extent, attributable to its incorporation of Lipschitz constraints. Subsequently, we further investigated the potential of ICLRNN by delving into the Long Range Arena benchmark [16], encompassing tasks involving long sequence modeling such as byte-level text sentiment classification and pixel sequence image classification. Both ICRNN and our ICLRNN exhibit limitations in tasks that require the modeling of very long time-dependent sequences, such as the byte-level text classification task, where both achieve an accuracy of approximately 50% in this binary sentiment classification task. This limitation stems from the challenge of the exploding gradient problem and lack of representative power due to non-negative weights, which both methods encounter when attempting to model extremely lengthy sequences, and we provide a theoretical explanation as follows.
@@
While the Lipschitz constraint may alleviate the exploding gradient problem by bounding the weights, integrating input convexity introduces a new challenge by raising this upper bound. Applying the Lipschitz constraint first to reduce the spectral norm to 1, followed by imposing the input convex constraint, yields a neural network no longer 1-Lipschitz constrained but rather $n\sqrt{r(W)}$-constrained, where $n$ represents the number of time-steps. Consequently, in modeling very long time-dependent sequences, the Input Convex structure exacerbates the exploding gradient problem, rendering it unmitigated even with Layer Normalization. For readers interested in tackling the challenges of the exploding and vanishing gradient problem within RNNs, we recommend exploring several noteworthy and comprehensive works [24, 25, 26]. We acknowledge this limitation as inherent to our proposed method, thereby constraining its applicability. However, despite this drawback, we opt to retain the Input Convex structure, as it demonstrates efficacy in many engineering applications where the length of time-series data is not as long as those in Long Range Arena benchmark. For example, in the first example where we develop a model for real-time optimization of CSTR, the model is to predict one sampling period only, since the controller is typically implemented in a feedback manner with the true measurements of states received at every time step. In the second case study where we develop a model for real-world solar irradiance prediction, again, the prediction horizon is relatively short (e.g., an hour or a few hours ahead) since this model is used for real-time energy planning. Therefore, there are many applications in engineering disciplines that focus on short-term prediction, where the proposed ICLRNN will play an important role due to its superiority in computational efficiency and robustness compared to traditional RNNs.
6 Conclusion
@@
This work introduces an Input Convex Lipschitz RNN to improve computational efficiency and non-adversarial robustness of neural networks in various engineering applications such as real-world solar irradiance prediction, and process modeling and optimization. The ICLRNN model was demonstrated to surpass state-of-the-art recurrent units across various engineering tasks, ranging from simulated studies to real-world applications, in both computational efficiency and non-adversarial robustness.
@@ Appendix A Proofs
Definition 1 ([27])
A function $f:X\rightarrow\mathbb{R}$ is convex if and only if the following inequality holds $\forall~{}(x_{1},x_{2})\in X$, with $x_{1}\neq x_{2}$, and $\forall~{}\lambda\in(0,1)$:
@@
Definition 2 ([28])
A function $f:X\rightarrow\mathbb{R}$ is Lipschitz continuous with Lipschitz constant $L$ (i.e., L-Lipschitz) if and only if the following inequality holds $\forall~{}(x_{1},x_{2})\in X$:
@@
The following lemmas are provided to prove Proposition 4 regarding the Lipschitz continuity of ICLRNN.
Lemma 1 ([28])
@@
$L(f)\leq L(g_{t})\times L(W_{t}^{y})\times L(g_{t-1})\times(L(W_{t-1}^{x})+L(W_{t-1}^{h})\times...L(g_{1})\times(L(W_{1}^{x})+L(W_{1}^{h})))$
If we ensure that the Lipschitz constants of all $W_{i}$ in ICLRNN are $\mathcal{O}(1)$, the ICLRNN for classification tasks with ReLU activation for hidden states and Softmax activation for output layer has a Lipschitz constant of $\mathcal{O}(1)$, while the ICLRNN for regression tasks with ReLU activation for hidden states and Linear activation for output layer has a Lipschitz constant of $\mathcal{O}(1)$ (i.e., Lemma 3 and Lemma 4).
The following lemmas are provided to prove Proposition 5 regarding the convexity of ICLRNN.
Lemma 5 ([27])
@@
Proof 2
By unrolling the recurrent operations over time $t$ in Eq. (1), the output of ICLRNN can be described as a function of Eq. (5). The output of an ICLRNN $y=f(x)$ is convex with respect to $x$ (i.e., Lemma 5 and Lemma 6).
@@
Appendix B MPC-based CSTR results
Table 2: Parameter values of CSTR.
$F=5~{}m^{3}/hr$ $V=1~{}m^{3}$ $R=8.314~{}kJ/kmol~{}K$
$T_{0}=300~{}K$ $C_{p}=0.231~{}kJ/kg~{}K$ $\rho_{L}=1000~{}kg/m^{3}$
$Q_{s}=0.0~{}kJ/hr$ $E=5\times 10^{4}~{}kJ/kmol$ $C_{A0_{s}}=4~{}kmol/m^{3}$
$k_{0}=8.46\times 10^{6}~{}m^{3}/kmol~{}hr$ $\Delta H=-1.15\times 10^{4}~{}kJ/kmol$
Table 1 indicates the parameter values for the CSTR. Fig. 1(a) showcases the model’s testing accuracy with respect to varying degrees of Gaussian noise applied to the input. These findings indicate ICLRNN’s non-adversarial robustness across diverse hypothesis space sizes with respect to different Gaussian noise levels. Notably, ICLRNN exhibits superior computational efficiency, as evidenced by its notably lower number of FLOPs compared to other models.
While ICRNN performs well against various degrees of Gaussian noise within smaller hypothesis spaces, it encounters challenges when the hypothesis space expands, succumbing to the issue of the exploding gradient problem. LRNN demonstrates comparable performance to ICLRNN specifically with a layer size of (512, 512). However, it is computationally demanding, characterized by substantially higher FLOPs (i.e., 9,992,212 for LRNN compared to 1,599,508 for ICLRNN). This computational intensity stands as a significant drawback despite its competitive performance.
We now delve into the control performance assessment of NN-based LMPC. To assess and compare the convergence rates of various NN-based LMPC, we conduct different NN-based LMPC under a fixed timeframe and repeat with different initial conditions. Figs. 3 and 4 provide compelling evidence that the ICLRNN-based LMPC achieves convergence, unlike other NN-based LMPC, within a fixed timeframe. This outcome distinctly indicates that ICLRNN-based LMPC exhibits the fastest convergence speed under identical settings.
In Fig. 3, the left panels show the results corresponding to the initial condition set at $[-1.5~{}kmol/m^{3},70~{}K]$. Specifically, the top left panel displays the concentration profile, the middle left panel shows the temperature, and the bottom left panel exhibits the converging path. Additionally, the right panels showcase the results from the initial condition at $[1.5~{}kmol/m^{3},-70~{}K]$, featuring the concentration profile in the top right panel, temperature profile in the middle right panel, and the converging path in the bottom right panel. Moving to Fig. 4, the left panels depict the results obtained from the initial condition of $[-1.25~{}kmol/m^{3},50~{}K]$. Here, the top left panel represents the concentration, the middle left panel illustrates the temperature, and the bottom left panel shows the converging path. Similarly, the right panels exhibit the results of the initial condition at $[1.25~{}kmol/m^{3},-50~{}K]$.
Figure 3: Concentration profile (top left), temperature profile (middle left), converging path (bottom left) in a fixed timeframe with an initial condition at $[-1.5~{}kmol/m^{3},70~{}K]$ under LMPC, and concentration profile (top right), temperature profile (middle right), converging path (bottom right) in a fixed timeframe with an initial condition at $[1.5~{}kmol/m^{3},-70~{}K]$.
Figure 4: Concentration profile (top left), temperature profile (middle left), converging path (bottom left) in a fixed timeframe with an initial condition at $[-1.25~{}kmol/m^{3},50~{}K]$, and concentration profile (top right), temperature profile (middle right), converging path (bottom right) in a fixed timeframe with an initial condition at $[1.25~{}kmol/m^{3},-50~{}K]$.
Appendix C Real-world solar irradiance prediction results
At present, the Solar PV system serves as the primary energy source for the industrial facility, while the main utility grid acts as a secondary energy source to supplement any deficiencies in solar energy production. Any surplus solar energy generated beyond the current requirements is efficiently stored in batteries for future utilization. Hence, in order to optimize real-time energy planning and mitigate manufacturing costs associated with energy procurement from the main utility grid, a high level of precision and accuracy in forecasting solar irradiance becomes imperative. Moreover, a simple and robust trained NN with relatively high accuracy can be seamlessly integrated into applications like energy dispatch optimization. To achieve these goals, the Solar Energy Research Institute of Singapore (SERIS) and Singapore Institute of Manufacturing Technology (SIMTech) have played pivotal roles by installing various sensors specifically designed for the Solar PV system. These sensors include irradiance sensors, humidity sensors, wind sensors, and module temperature sensors. Drawing on parameters akin to those employed by [23], we utilize minute-based average global solar irradiance, ambient humidity, module temperature, wind speed, and wind direction to predict solar irradiance in the subsequent minute.
@@ Appendix C Real-world solar irradiance prediction results
Fig. 5 illustrates the production pipeline adopted by LHT Holdings. Meanwhile, in Fig. 6(a), a depiction of the different energy sources employed within the production process is presented. Fig. 6(b) showcases the physical setup of the Solar PV system, encompassing solar panels, inverters, batteries, meters, humidity and wind sensors, module temperature sensors, and irradiance sensors. Data is systematically recorded on a minute-by-minute basis and seamlessly uploaded online into the system.
@@
(a) LHT Holdings energy sources for manufacturing.
(b) LHT Holdings Solar PV system.
Figure 6: Performance of solar irradiance forecasting.
@@
Moving to the bottom two figures, the LRNN and ICLRNN prediction results for the same date, December 28, 2023, are presented. The LRNN, constrained by Lipschitz continuity, exhibits significantly improved performance compared to the RNN and LSTM models. However, it still encounters challenges in accurately predicting solar irradiance during sudden fluctuations, and its excessively high FLOPs signify its computational intensity (i.e., 2,498,562 for LRNN compared to 399,362 for ICLRNN). Moreover, slight fluctuations in the LRNN predictions occur during periods of minimal solar irradiance at the start and end of the day. Notably, these observations persist across all trials and are not specific solely to December 28, 2023. It is clearly shown in Fig. 2(a) that our ICLRNN model surpasses state-of-the-art models, demonstrating superior predictive capabilities across various scenarios.
@@
Appendix D Computing resources
The experiments were conducted on the NVIDIA Quadro RTX A5000 GPU with 24 GB memory, but they are also compatible with CPUs equipped with at least 64 GB of RAM. A single run to train the real-world solar irradiance prediction model took about 6 hours while a single run for NN-MPC is about 20 minutes using CPU with 64 GB of RAM.